OpenStudy (anonymous):

Match each each equation equation with the correct solution. 2e^x - 1 = 0 Choose... -1.2345/ 0.2877 /0.2747 /0.6438/ -0.06931/ 1.5041

OpenStudy (anonymous):

ok lets solve this can be written as 2e^x=1

OpenStudy (anonymous):

take ln (natural log) of both sides ln(2e^x)=ln(1)

OpenStudy (anonymous):

simplify the above let me know what you get after this.

OpenStudy (anonymous):

@brendas

OpenStudy (anonymous):

write 2e^x-1=0 as e^x=1/2 take log of both sites ln(e^x)=ln(1/2) using property of log xln(e)=ln(1/2) since ln(e)=1 x=ln(1/2) now use your calculator

OpenStudy (anonymous):

@brendas if you need help about this let me know.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

ok what you are unable to fallow in the above?

OpenStudy (anonymous):

i dont get how to solve it

OpenStudy (anonymous):

ok can i write 2e^x-1=0 as 2e^x=1 reply

OpenStudy (anonymous):

yes but where does the 0 go

OpenStudy (anonymous):

opps:P

OpenStudy (anonymous):

let me do it again 2e^x-1=0 add one to the both sides of equation 2e^x-1+1=0+1 2e^x=0+1 2e^x=1 (0+1 is equal to 1)

OpenStudy (anonymous):

is it ok now ?

OpenStudy (anonymous):

@brendas

OpenStudy (anonymous):

yea better

OpenStudy (anonymous):

ok so whats the next step

OpenStudy (anonymous):

ok now divide both sides by 2 \[\frac{2e^x}{2}=\frac{1}{2}\] since 2/2=1 so \[e^x=\frac{1}{2}\] are you with me ? are you getting these steps?

OpenStudy (anonymous):

yea i get the step

OpenStudy (anonymous):

ok now take natural log of both sides \[\ln(e^x)=\ln(\frac{1}{2})\] using property of log the the thing in exponents comes down and gets multiplied with log so x should come down \[xln(e)=\ln(\frac{1}{2})\] is it ok ?

OpenStudy (anonymous):

am here im looking at the next step which i dont get now

OpenStudy (anonymous):

it is the property of log that the exponent comes down and gets multiplied like \[\log_{b}x^n=nlog_{b}x\] is it ok now

OpenStudy (anonymous):

im sorry but i just dont get the log parts!!!!! so whats my next step on the problem

OpenStudy (anonymous):

this is the answer i get -2.3026

OpenStudy (anonymous):

ok take a look on in your book about logs section. we were at \[xln(e)=\ln \frac{1}{2}\] since ln(e)=1 \[x=\ln (\frac{1}{2})\] got it ?

OpenStudy (anonymous):

got a calculator?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok use it and find ln(1/2) let me know what you got?

OpenStudy (anonymous):

@brendas you got it?

OpenStudy (anonymous):

ok i found using calculator \[x=-0.6931\] this the solution Best of Luck with rest of your studies.